Please answer
You are swimming underwater in a large lake when you find a rock and decide to throw it down (all while fully immersed in the water). If one assumes that the relationship between fluid resistance and speed is given by this equation: f-kv, then compute the speed of the rock as a function of time as it travels down in the water given that you throw it with a speed of 3mg/k. (Note that k is the coefficient in the above fluid resistance equation.) Express your answer in terms of vr (terminal speed), k, m, and t.
Answer #1

We know that the terminal speed of the rock will be when net force will be Zero.
Therefore
f- mg = 0
kv = mg
at this point velocity would be terminal therefore , v = vT
kvT = mg
vT = mg/k ------------(1)
Now at any point
mg - kv = ma
we know that acceleration is given by
a= dv/dt , therfore
mg - kv = m*(dv/dt)
Now ,
du = dt
Now integrating both side , we get
7n dv U) dt 7n
-n
where C is integration constant
-m
taking anti-log both side
mq-ku
kv =mg - e^{left ( t+C ight )left ( rac{-m}{k} ight )}
Now at t = 0 , v = 3mg/K
kleft ( rac{3mg}{k} ight ) =mg - e^{left ( 0+C ight )left ( rac{-m}{k} ight )}
3mg =mg - e^{left ( C ight )left ( rac{-m}{k} ight )}
2mg = e(FTC)
on taking log both side
nzmg
C = {left ( rac{k}{m*ln(2mg)} ight )}
Therefore the equation of v will be
kv =mg - e^{left ( t+C ight )left ( rac{-m}{k} ight )}
v =rac{mg}{k} - left ( rac{1}{k} ight )e^{left ( t+C ight )left ( rac{-m}{k} ight )}
using the equation of vT
v =v_{ au } - left ( rac{1}{k} ight )e^{-left ( rac{mt}{k} + rac{1}{ln(2mg)} ight )}
KV ng